A train travels along a horizontal circular curve that has a radius of 200 m. If the speed of the train is uniformly increased from 30 km/h to 45 km/h in 5 s, determine the magnitude of the acceleration at the instant the speed of the train is 40 km/h.

Grade:Upto college level

1 Answers

Yogita Bang
11 years ago

a = aT + aR ----------->(vector addition)

Therefore, a2= aT2 + aR   (since, angle between aT & aR is 90º)

aT= {45(5/18) - 30(5/18)}/5 = 5/6

aR= v2/r = {40(5/18)}/200 = 1/18

a2= 25/36 + 1/324 = 226/324

a≈ 0.835 m/s2